Company has a contract to sell 10,000 parts. Each part will sell for $45.50 apiece. There is a 20% probability that expeses will be $40,000 and an 80% probability that expenses will be $50,000. What is the expected pretax income for the firm's sales?
Answer by edjones(8007)

10000*45.50=455,000 gross income
.2*40000 + .8*50000=48,000
455000-48000=$407,000.
.
Ed